Sabrina lived half her life in Toronto and one third of her life in Niagara Falls. She then moved to Brampton.
1) If She lived in Brampton for four years, how old is Sabrina now?

2) How long has she lived in each place

Answer by richwmiller(17219)

x/2+x/3+4=x
3x+2x+24=6x
5x+24=6x
24=x
12 years in Toronto
8 years in NF
and 4 years in Brampton.